# Learning not cheating: AI assistance can enhance rather than hinder skill development

### v2

@@ Appendix C Results Study 3
| Condition | Main Sample | Followup Sample |
| Practice wo AI | ||
| Practice w AI | ||
@@
| Overall | Practice wo AI | Practice w AI | See AI example | p | SMD | |
@@
C2 AI examples improve writing skill
@@
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Followup | LM | ER | EN | F | ER | |
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
C3 Seeing AI examples was less effortful
@@
| Means — (SE) | |||||
| Practice wo AI | 2.83 | 5.01 | 6.17 | 9.00 | 259.34 |
| (.270) | (.541) | (.642) | (1.495) | (99.161) | |
| Practice w AI | 2.71 | 4.05 | 5.89 | 8.65 | 228.45 |
| (.270) | (.540) | (.641) | (1.493) | (99.050) | |
| See AI example | 1.85 | .81 | 5.52 | 4.99 | 24.98 |
| (.271) | (.542) | (.643) | (1.499) | (99.392) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | -.14* | -.55*** | -.14* | -.07 | -.10 |
| (.056) | (.056) | (.057) | (.056) | (.056) | |
| Practice wo AI vs. See AI example | -1.13*** | -2.41*** | -.32*** | -.83*** | -.73*** |
| (.059) | (.067) | (.056) | (.057) | (.056) | |
| Practice w AI vs. See AI example | -.99*** | -1.86*** | -.18** | -.76*** | -.64*** |
| (.058) | (.063) | (.056) | (.057) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
@@
| Means — (SE) | |||||
| Practice wo AI | 2.45 | 5.40 | 7.05 | 6.14 | 432.01 |
| (.170) | (.562) | (.623) | (.655) | (109.856) | |
| Practice w AI | 2.50 | 5.52 | 7.19 | 6.35 | 466.90 |
| (.169) | (.561) | (.622) | (.654) | (109.733) | |
| See AI example | 2.57 | 5.86 | 7.21 | 6.63 | 517.26 |
| (.170) | (.563) | (.625) | (.656) | (110.112) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | .09 | .06 | .07 | .10 | .10 |
| (.057) | (.056) | (.057) | (.057) | (.056) | |
| Practice wo AI vs. See AI example | .22*** | .25*** | .08 | .23*** | .24*** |
| (.056) | (.055) | (.056) | (.056) | (.055) | |
| Practice w AI vs. See AI example | .13* | .19*** | .01 | .13* | .14* |
| (.056) | (.056) | (.056) | (.056) | (.055) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
@@
| Means — (SE) | |||||
| Practice wo AI | .20 | .15 | .18 | .31 | .34 |
| (.155) | (.272) | (.085) | (.278) | (.167) | |
| Practice w AI | .30 | .22 | .25 | .45 | .49 |
| (.155) | (.272) | (.085) | (.277) | (.167) | |
| See AI example | .51 | .97 | .28 | 1.08 | .62 |
| (.155) | (.273) | (.085) | (.278) | (.168) | |
| Effect Sizes (d) — (SE) | |||||
| Practice wo AI vs. Practice w AI | .21*** | .08 | .25*** | .15** | .27*** |
| (.057) | (.057) | (.057) | (.057) | (.057) | |
| Practice wo AI vs. See AI example | .62*** | .94*** | .38*** | .86*** | .53*** |
| (.057) | (.058) | (.056) | (.058) | (.057) | |
| Practice w AI vs. See AI example | .41*** | .86*** | .13* | .71*** | .25*** |
| (.057) | (.058) | (.056) | (.057) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | |||||
C4 Seeing an AI example did not discourage motivation for future learning
@@
| extbfEffect Sizes (d) | |||
| Practice wo AI vs. Practice w AI | -.01 | .02 | 1.19 |
@@
C5 The effects of seeing an AI example persist
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | |||||||||
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disagregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||
C6 Seeing AI examples was equally effective across subgroups

### v3

@@ Appendix F Results Study 5
| Condition | Main Sample | Follow-up Sample |
| Practice w/o AI | ||
| Practice w/ AI | ||
@@
| Overall | Practice w/o AI | Practice w/ AI | See AI example | p | SMD | |
@@
F2 AI examples improve writing skill
@@
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disaggregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
| GPT-4o | Claude | Ex. Controls | Ex. Cheaters | Follow-up | LM | ER | EN | F | ER | |
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disaggregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||||||||
F3 Seeing AI examples was less effortful
@@
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
| Means | ||||||
| Practice wo AI | 2.83 | 5.01 | 6.17 | 9.00 | 259.34 | 27.62 |
| (.270) | (.541) | (.642) | (1.495) | (99.161) | (13.213) | |
| Practice w AI | 2.71 | 4.05 | 5.89 | 8.65 | 228.45 | 21.20 |
| (.270) | (.540) | (.641) | (1.493) | (99.050) | (13.200) | |
| See AI example | 1.85 | .81 | 5.52 | 4.99 | 24.98 | -8.17 |
| (.271) | (.542) | (.643) | (1.499) | (99.392) | (13.246) | |
| Effect Sizes (d) | ||||||
| Practice wo AI vs. Practice w AI | -.14* | -.55*** | -.14* | -.07 | -.10 | -.15** |
| (.056) | (.056) | (.057) | (.056) | (.056) | (.056) | |
| Practice wo AI vs. See AI example | -1.13*** | -2.41*** | -.32*** | -.83*** | -.73*** | -.84*** |
| (.059) | (.067) | (.056) | (.057) | (.056) | (.056) | |
| Practice w AI vs. See AI example | -.99*** | -1.86*** | -.18** | -.76*** | -.64*** | -.69*** |
| (.058) | (.063) | (.056) | (.057) | (.056) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
@@
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
| Means | ||||||
| Practice wo AI | 2.45 | 5.40 | 7.05 | 6.14 | 432.01 | 68.46 |
| (.170) | (.562) | (.623) | (.655) | (109.856) | (19.117) | |
| Practice w AI | 2.50 | 5.52 | 7.19 | 6.35 | 466.90 | 72.19 |
| (.169) | (.561) | (.622) | (.654) | (109.733) | (19.090) | |
| See AI example | 2.57 | 5.86 | 7.21 | 6.63 | 517.26 | 77.84 |
| (.170) | (.563) | (.625) | (.656) | (110.112) | (19.159) | |
| Effect Sizes (d) | ||||||
| Practice wo AI vs. Practice w AI | .09 | .06 | .07 | .10 | .10 | .06 |
| (.057) | (.056) | (.057) | (.057) | (.056) | (.057) | |
| Practice wo AI vs. See AI example | .22*** | .25*** | .08 | .23*** | .24*** | .15** |
| (.056) | (.055) | (.056) | (.056) | (.055) | (.056) | |
| Practice w AI vs. See AI example | .13* | .19*** | .01 | .13* | .14* | .09 |
| (.056) | (.056) | (.056) | (.056) | (.055) | (.056) | |
| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
@@
| sqrt(Time) | log(Keystrokes) | Subjective Rating (0 - 10) | Time | Keystrokes | Keystrokes/min | |
| Means | ||||||
| Practice wo AI | .20 | .15 | .18 | .31 | .34 | .20 |
| (.155) | (.272) | (.085) | (.278) | (.167) | (.277) | |
| Practice w AI | .30 | .22 | .25 | .45 | .49 | .30 |
| (.155) | (.272) | (.085) | (.277) | (.167) | (.277) | |
| See AI example | .51 | .97 | .28 | 1.08 | .62 | 1.00 |
| (.155) | (.273) | (.085) | (.278) | (.168) | (.278) | |
| Effect Sizes (d) | ||||||
| Practice wo AI vs. Practice w AI | .21*** | .08 | .25*** | .15** | .27*** | .12* |
| (.057) | (.057) | (.057) | (.057) | (.057) | (.057) | |
| Practice wo AI vs. See AI example | .62*** | .94*** | .38*** | .86*** | .53*** | .90*** |
| (.057) | (.058) | (.056) | (.058) | (.057) | (.058) | |
| Practice w AI vs. See AI example | .41*** | .86*** | .13* | .71*** | .25*** | .78*** |
| (.057) | (.058) | (.056) | (.057) | (.056) | (.058) | |
| Note. *** p < .001, ** p < .01, * p < .05. | ||||||
F4 Seeing an AI example did not discourage motivation for future learning
@@
| Effect Sizes (d) | |||
| Practice w/o AI vs. Practice w/ AI | -.01 | .02 | 1.19 |
@@
F5 The benefits of seeing an AI example were just as large a day later
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disaggregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | |||||||||
@@
| Note. GPT-4o is the main specification. Ex. Controls is the main specification, unadjusted for demographic and pretreatment variables, Ex. Cheaters excludes the 3% of participants who admitted to cheating on the test phase. LM to ER are disaggregated scores for each of the five principles. LM = Less is More, ER = Easy Reading, EN = Easy Navigation, F = Formatting, ER = Easy Responding. *** p < .001, ** p < .01, * p < .05. | ||||
As mentioned in the main text the effect of practicing with AI and seeing an AI example did not become attenuated one day later. See Table S34.
| Term | Estimate | SE | t | p-value |
| Intercept | -8.469 | 4.502 | -1.881 | 0.060 |
| Condition: Practice w AI | 0.421 | 0.076 | 5.551 | 0.000 |
| Condition: See AI example | 0.467 | 0.075 | 6.253 | 0.000 |
| Condition: Practice w/ AI Follow-up | -0.007 | 0.158 | -0.044 | 0.965 |
| Condition: See AI example Follow-up | -0.006 | 0.157 | -0.036 | 0.971 |
| Phase: Follow-Up | 0.356 | 0.114 | 3.139 | 0.002 |
| Pretest score | 0.394 | 0.031 | 12.877 | 0.000 |
| Year of birth | 0.006 | 0.002 | 2.524 | 0.012 |
| Gender: Male | -0.096 | 0.055 | -1.724 | 0.085 |
| Gender: Other | 0.117 | 0.249 | 0.470 | 0.639 |
| Race/ethnicity: White | 0.215 | 0.117 | 1.846 | 0.065 |
| Race/ethnicity: Black | -0.140 | 0.125 | -1.125 | 0.261 |
| Race/ethnicity: Asian | 0.575 | 0.148 | 3.880 | 0.000 |
| Race/ethnicity: Latino | 0.216 | 0.123 | 1.756 | 0.079 |
| Race/ethnicity: Other | -0.180 | 0.250 | -0.719 | 0.472 |
| Education: High school graduate (high school diploma or equivalent including GED) | -0.184 | 0.397 | -0.464 | 0.642 |
| Education: Some college but no degree | -0.187 | 0.394 | -0.475 | 0.635 |
| Education: Associate degree in college (2-year) | -0.207 | 0.399 | -0.518 | 0.605 |
| Education: Bachelor’s degree in college (4-year) | -0.121 | 0.390 | -0.310 | 0.756 |
| Education: Master’s degree | -0.033 | 0.393 | -0.085 | 0.933 |
| Education: Doctoral degree (PhD) | -0.227 | 0.439 | -0.518 | 0.604 |
| Education: Non-PhD Professional degree (JD, MD) | 0.099 | 0.456 | 0.218 | 0.828 |
| Writing skill | 0.045 | 0.017 | 2.578 | 0.010 |
| Motivation: Hardly motivated | -0.196 | 0.258 | -0.760 | 0.447 |
| Motivation: Somewhat motivated | -0.142 | 0.243 | -0.582 | 0.561 |
| Motivation: Very motivated | -0.338 | 0.244 | -1.388 | 0.165 |
| Motivation: Extremely motivated | -0.430 | 0.249 | -1.731 | 0.084 |
| Experience: I have tried AI writing assistant(s) but hardly ever use them | 0.132 | 0.079 | 1.678 | 0.093 |
| Experience: I use AI writing assistant(s) a few times per week | 0.003 | 0.094 | 0.034 | 0.973 |
| Experience: I use AI writing assistant(s) about once a week | 0.070 | 0.095 | 0.730 | 0.465 |
| Experience: I use AI writing assistant(s) every day | -0.063 | 0.131 | -0.480 | 0.631 |
F6 Seeing AI examples was equally effective across subgroups
@@
F7 Pairwise Comparisons
The relative likelihood of a cover letter receiving an invitation to an interview was correlated with the GPT-rated writing quality. See Figure S11.
As shown in Table S36, participants who had practiced writing cover letters with AI were more likely to be invited to a hypothetical job interview.
| Without controls | With controls | |
| Means (SE) | ||
| Practice w AI | .50 | .51 |
| (.01) | (.01) | |
| Practice wo AI | .48 | .48 |
| (.01) | (.01) | |
| See AI example | .51 | .51 |
| (.01) | (.01) | |
| Model coefficients (SE) | ||
| Precision () | 16.17*** | 7.77*** |
| (1.47) | (.51) | |
| Symmetry (Log()) | .35*** | -.23*** |
| (.08) | (.07) | |
| Pairwise comparison (SE) | ||
| Practice w AI - Practice wo AI | .08** | .09* |
| (.03) | (.04) | |
| Practice w AI - See AI example | -.02 | -.03 |
| (.03) | (.04) | |
| Practice wo AI - See AI example | -.10*** | -.12** |
| (.03) | (.04) | |
| Statistics | ||
| N | 1934.00 | 1916.00 |
| AIC | 2447.07 | 2518.06 |
| BIC | 2474.91 | 2684.80 |
| log(Likelihood) | -1218.54 | -1229.03 |